# Black Hole from Entropy Maximization

### v4

@@ VI.2.1 Meaning of the upper bound (19)
We here have $|-\tilde{s}^{\mu}k_{\mu}|=|-\tilde{s}^{t}k_{t}|=\tilde{s}^{t}(-g_{tt})|k^{t}|$. Also, for excitations consistent with thermodynamics, $\langle-T^{t}{}_{t}\rangle\sim\langle T^{r}{}_{r}\rangle$ holds Landau_SM , and (VI.2) means $\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu}\sim g_{rr}|k^{r}|^{2}\langle-T^{t}{}_{t}\rangle$. Using these, (52) leads to
@@
We discuss the meaning of this $\Delta\zeta$. In general, when considering a light sheet going to a horizon, the light sheet must not be continued to the interior, since the horizon entropy contains the contribution from objects that fell inside: Otherwise, it would be counted twice Bousso2 . In the geometry (7) without trapped surface, the “would-be horizon” for a point $r$ exists at $r=a(r)$. Therefore, our $\Delta\zeta$ represents the affine distance from a point to the “would-be end point”, and hence our $\Delta\zeta$ corresponds to that of Eq.(1.9) in Ref.FMW .
@@ VI.2.2 Exact saturation of a local sufficient condition by $g_{\mu\nu}^{*}$
The interior metric (46) saturates (57) exactly at the leading order for $r\gg l_{p}$: The left hand side becomes $\frac{\hbar\alpha_{*}}{2\pi}\frac{s_{*}}{4\pi r^{2}}-\frac{\hbar}{2\pi}\frac{1}{\sqrt{g^{*}_{rr}}}\partial_{r}\left(\frac{s_{*}}{4\pi r^{2}}\right)\approx\frac{\hbar\alpha_{*}}{2\pi}\frac{s_{*}}{4\pi r^{2}}=\frac{1}{4\pi G\eta r^{2}}$ for $r\gg l_{p}$ (from (39) and (41)), while the right one becomes $\langle-T^{t}{}_{t}\rangle_{*}+\langle T^{r}{}_{r}\rangle_{*}=\frac{1}{4\pi G\eta r^{2}}$ (from (IV.3)). Thus, the gravity condensate saturates not only the global bound (5) but also the local condition (54) relevant to the local structure. This is so non-trivial that there must be something behind it.
@@
The identity of a black hole is still unknown, and the microscopic constituents should involve fundamental degrees of freedom in quantum gravity, where there is no notion of classical spacetime geometry. Motivated by thermodynamics and (local) holography, in the present paper, we adopted the characterization that a black hole maximizes thermodynamic entropy for a given surface area. For spherical static configurations, we uniquely obtained the entropy-maximizing configuration consistent with local thermodynamics and the 4D semi-classical Einstein equation with many matter fields. That is, a self-gravitating collection of near-Planckian excited quanta condensates into the radially-uniform dense configuration without horizons or singularities as in Fig.2, where the self-gravity and large quantum pressure are balanced (called semi-classical gravity condensate). The interior metric (46) (say, for $n_{s}$ scalar fields, with the parameters (33)) satisfies the semi-classical Einstein equation self-consistently and non-perturbatively in $\hbar$. The bulk dynamics makes the entropy of the interior quanta exactly follow the Bekenstein-Hawking formula, leading to the Bousso bound for thermodynamic entropy and the saturation of the local sufficient conditions. This is a candidate picture of black holes in quantum theory.
Let us now discuss 10 future prospects and speculate on what possibilities there are for a more complete quantum description of black holes and a construction of quantum gravity.
@@
A model is in the framework of group field theory, a second quantization of loop quantum gravity Oriti1 . Gluing spherically-symmetric quantum-gravity states kinematically and maximizing the entropy for a given surface area, a quantum gravitational configuration appears with the entropy proportional to the surface area, and the coefficient is fixed by using the Unruh effect and thermodynamic relations. A remarkable point is that the holographic property of the entropy of the quanta living in the interior bulk holds for any size $r$, which corresponds to the radial uniformity in our semi-classical gravity condensate. (See Sec.8 of Ref.Y1 for more discussions.)
@@
Then, interesting similarities can be observed between $n$ and $\mathcal{N}$. For our condensate, the temperature $T_{loc}(r)\sim\frac{m_{p}}{\sqrt{n}}$, characteristic wavelength $\lambda_{*}(r)\sim\sqrt{n}l_{p}$, and energy $\Delta E_{loc}^{*}(r)\sim\sqrt{n}m_{p}$ are controlled by $n$. These manifestations of $n$ are exactly the same as those of $\mathcal{N}$ in the condensate of Ref.Dvali3 , although the former is for local subsystems and the latter for the entire system. Also, from (59), the condition $n\gg 1$ corresponds to that for the classicality in Ref.Dvali3 . Therefore, a possibility is that realizing the condensate in Ref.Dvali3 in local subsystems and connecting them consistently with the semi-classical Einstein equation lead to our gravity condensate.
Note that we have $N(r)\neq N_{g}(r)$ for a generic configuration, but our gravity condensate satisfies
@@
The two other models above discuss quantum gravitational effects without considering quantum matter ones, while our model describes matter quanta in the self-consistent (and non-perturbative) classical gravitational field. Therefore, a more detailed study of the connection between the three models would provide some clues to the above expectation.
@@
For the gravity condensate, (61) is saturated: $\lambda_{*}(r)\sim\mathcal{R}_{*}(r)^{-1/2}$ (from (35)), and also the equivalence in the number of gravity and matter quanta (60) holds. Thus, the correspondence (62) would imply that, the gravity condensate phase is governed essentially by quantum-gravitational effects while due to the large occupation number (60) for $n\gg 1$, the mean-field approximation of quantum gravity by (1) Kiefer holds albeit barely, leading to the semi-classical description by the metric (46).
A way to study this speculation is to first find an appropriate order parameter, study a Ginzburg-Landau-like theory or Gross–Pitaevskii-like equation including the effect of the redshift, and reproduce the radial uniformity or the interior metric. One could then consider the quantum many-body model behind it.
We finally discuss the notion of equilibrium in a self-gravitating system. In the absence of self-gravity, the most typical configuration maximizes the entropy for given macroscopic parameters and corresponds to the equilibrium state. Therefore, the derivation of the gravity condensate as the entropy-maximizing configuration suggests that the gravity condensate is the self-gravitating equilibrium state for a given surface area. This is consistent with the result of Appendix D.
Here, the effect of interactions should be important. In general, a phase is a uniform equilibrium state which is achieved through internal interactions Landau_SM . The parameter $\eta$ was originally introduced as a phenomenological parameter to represent the scattering/interaction effects of Hawking-like radiation by the gravitational potential and matters inside; a larger $\eta$ represents a larger interaction, whose macroscopic effect appears in the radial pressure $\langle T^{r}{}_{r}(r)\rangle$ in (IV.3) KY1 . Therefore, the above quantum many-body model should have interactions between matter and gravity quanta, which should explain $\eta$ and the large redshift $g_{tt}(r)\sim-e^{-\frac{R^{2}-r^{2}}{2\eta\sigma}}$ (or the time scale $\Delta t\sim\mathcal{O}(e^{a_{0}^{2}/l_{p}^{2}})$ in Sec.V.1) microscopically.
@@
5. Thermodynamic entropy vs entanglement entropy—. Another field-theoretic approach to $S[g_{\mu\nu};R)$ would be to formulate it as entanglement entropy. In this case, the Unruh temperature is often applied to local Rindler regions as the local temperature Minic ; Casini ; Jacobson_entangle , but it does not necessarily agree with the thermodynamic temperature consistent with the self-consistent gravity. In the case of, say, self-gravitating thermal radiation, the Unruh temperature is different from the local temperature obtained by applying Stefan-Boltzmann law or Tolman’s law in the metric (see Appendix A). This is natural because the local temperature of an object in global thermodynamic equilibrium, which is fixed by Tolman’s law in the self-consistent metric, does not generically coincide with the Unruh temperature determined by the acceleration required to stay against the self-gravity. Therefore, the entanglement entropy calculated by such methods can be different from the thermodynamic entropy $S[g_{\mu\nu};R)$. The crucial point is whether to take into account the self-gravity determined by the semi-classical Einstein equation.
As argued in Sec.V.1.1, however, the local temperature (39) of the gravity condensate can be obtained both kinematically (from the Unruh formula) and dynamically (due to the particle creation). This is based on the self-consistent metric (46), where the relation $\lambda_{*}(r)\sim\mathcal{R}_{*}^{-\frac{1}{2}}(r)$ holds (possibly involving quantum gravity effects as mentioned in Prospect 3). Thus, we expect that applying such field-theoretic techniques locally enables us to evaluate the entropy density of the gravity condensate as entanglement entropy, and the entanglement entropy and the thermodynamic (Boltzmann) entropy agree and give the Bekenstein-Hawking formula. Indeed, this holds for the quantum gravity condensate proposed in Ref.Oriti1 . We would like to check this expectation in the future.
@@
Furthermore, the origin of the Bekenstein-Hawking entropy should be consistent with a physical understanding of information problem (see below): through the unitary evolution, the microscopic states contributing to the entropy should correspond one-to-one to the initial states of a collapsing matter forming the black hole.
@@
6. Physical resolution of the information problem—. The purification of the initial Minkowski vacuum after evaporation and the “formal” derivation of the Page curve are only part of the information problem. More essentially, we need to understand the dynamical mechanism by which the initial wavefunction of a collapsing matter forming the black hole is recovered after evaporation, leading to the physical Page curve as the result of the unitary evolution of matter and gravity. We must also clarify the interior structure with quantum dynamics that resolves the singularity. Such overall consistency will show the true identity of the black hole.
@@
In the semi-classical gravity condensate, the information of a (typical) collapsing matter is stored in the bulk interior, and the amount agrees with the Bekenstein-Hawking formula. The condensate has neither horizon nor singularity (although we still need to understand the small center part at a fully quantum-gravity level), and almost all parts evaporate in the vacuum due to the Hawking-like radiation. Therefore, it should be natural to expect that (most of) the information recovers after the evaporation. As mentioned above, however, we still need to clarify the mechanism consistent with the energy flow: how the information of the matter leaks out gradually during the evaporation. A possibility is that scattering between a collapsing matter and the radiation occurs frequently at each point inside KY2 , and such interactions should transfer the information to the emitted radiation, reproducing the physical Page curve.
@@
To see this explicitly, one would at least need to check the initial-state dependence of the radiation emitted after the interaction. However, the condensate is in a typical state, and such dependence is not easy to see. One idea to overcome this is to consider perturbations from the typical state, adding a small atypical portion to the condensate and tracking it during evaporation. Another is to study how to distinguish between two typical states in a certain protocol and implement it in this model.
@@
7. Non-typical black holes—. The semi-classical gravity condensate has been obtained as the entropy-maximizing configuration for a given surface area and thus should be the most typical black hole. As described around Fig.4, it can be formed by adiabatic processes in a heat bath of Hawking temperature. Then, what is a non-typical black hole that is formed by a generic collapse of matter not by such a reversible process?
@@
A way to consider it is to analyze the time evolution of a collapsing matter including the backreaction from particle creation during the collapse (see footnote 15). Indeed, we can consider the matter as consisting of many spherical shells with small energy, distributed depending on the initial configuration and solve (1) self-consistently KY2 . It shows that the dense structure with the metric (46) is formed only around the surface, while the structure in deeper regions depends on the details of the initial distribution. The entropy is smaller than (42), and this is a non-typical black hole.
@@
Here, it is important to consider a finite width even when discussing collapse of a spherical shell. For simplicity it is often modeled by an infinitely thin shell BD ; BHmodel , but physically, any collapsing matter is an excitation of quantum fields with a physical information $|\psi\rangle$, which cannot be localized completely in the radial direction. Considering a finite-width shell as a collection of many tiny shells and including the evaporation effect, we can see that the tangential pressure $\langle T^{\theta}{}_{\theta}\rangle$ occurs dynamically and the gravity-condensate structure appears only around $r=a_{0}$ while the interior is vacuum KY2 ; KY4 . This is a result of the 4D dynamics; in the 4D conservation law $\nabla_{\mu}\langle T^{\mu}{}_{\nu}\rangle=0$, which contains $\langle T^{\theta}{}_{\theta}\rangle$, it can be large due to the 4D Weyl anomaly KY3 , while in the 2D cases, there is no tangential direction.
@@
Another way is to consider a solution space, consistent with the 4D Weyl anomaly, of the semi-classical Einstein equation and study various configurations in a non-perturbative manner for $\hbar$ HKLY . It shows that the most typical ones for a given surface area are similar to the semi-classical gravity condensate, and the number of such configurations explains the Bekenstein-Hawking entropy (except for $O(1)$ numerical factor). Also, non-typical ones have the dense structure only around the surface, which is consistent with the one above.
@@
Therefore, these different studies based on non-perturbative dynamics of the semi-classical Einstein equation give almost the same picture of the typical and non-typical black holes. It is interesting to study a time-dependent perturbation and consider how such a non-typical one evolves to the typical one.
@@
8. Relation to the classical picture of black holes—. One might wonder why the gravity condensate is so different from the classical picture of black holes, a vacuum region surrounded by a horizon. First, the classical one was originally derived from the classical dynamics of a collapsing matter and has not been confirmed observationally yet. Second, even at the semi-classical level, infinitely thin-shell models with 2D approximation are often used, leading to almost the same picture as the classical one BHmodel . In general, however, quantum fluctuations of modes with arbitrary angular momentum are induced in 4D spherically symmetric spacetime (which leads to 4D Weyl anomaly, for example). As mentioned above, the consideration of a finite width is important in the information problem. Furthermore, the backreaction from evaporation during the collapse is often neglected, but must be included (see footnote 15). On the other hand, our argument takes these points into account and obtains the picture of the gravity condensate. Thus, the difference appears.
@@
Another aspect in the difference between the conventional picture and the gravity condensate is that they belong to different branches in a solution space of the semi-classical Einstein equation. In Ref.HKLY , we constructed a self-consistent equation of the energy distribution $\frac{a(r)}{2G}$ including the effect of the 4D Weyl anomaly, and examined the structure of the solution space in a non-perturbative manner for $\hbar$. We then found that there exist two branches: one is perturbative and contains Schwarzschild-like metrics, while the other is non-perturbative and includes the dense solution (32). The point is that, the higher derivative term $\Box R_{RicciScalar}$, which appears generically in the anomaly BD , causes transitions between the two branches as $r$ changes. Therefore, it would be interesting to study 4D semi-classical time evolution of a collapsing matter including such higher derivative effects and see how the gravity condensate is generically formed.
@@
9. Gravitational field with finite entropy—. We briefly comment on entropy in gravity. An interesting aspect in estimating the entropy $S[g_{\mu\nu};R)$ in Sec.II is that local typicality and Hamiltonian constraint assign a finite entropy to a gravitational field $g_{\mu\nu}$. This is reminiscent of two things. First, it is similar to the idea that the Einstein equation corresponds to the equation of state in spacetime thermodynamics Jacobson ; Pad . Second, it should be relevant to the view of entropy as a gravitational charge Wald ; SY (see Sec.8 of Ref.Y1 for details). These could be relevant each other.

### v5

@@ VI.2.1 Meaning of the upper bound (19)
We here have $|-\tilde{s}^{\mu}k_{\mu}|=|-\tilde{s}^{t}k_{t}|=\tilde{s}^{t}(-g_{tt})|k^{t}|$. Also, for excitations consistent with thermodynamics, $\langle-T^{t}{}_{t}\rangle\gtrsim\langle T^{r}{}_{r}\rangle$ holds Landau_SM , and (VI.2) means $\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu}\sim g_{rr}|k^{r}|^{2}\langle-T^{t}{}_{t}\rangle$. Using these, (53) leads to
@@
We discuss the meaning of this $\Delta\zeta$. In general, when considering a light sheet going to a horizon, the sheet must not be continued to the interior, since the horizon entropy contains the contribution from objects that fell inside: otherwise, it would be counted twice Bousso2 . In the geometry (7) without trapped surface, the “would-be horizon” for a point $r$ exists at $r=a(r)$. Therefore, our $\Delta\zeta$ represents the affine distance from a point to the “would-be end point”, and hence our $\Delta\zeta$ corresponds to that of Eq.(1.9) in Ref.FMW .
@@ VI.2.2 Exact saturation of a local sufficient condition by $g_{\mu\nu}^{*}$
The interior metric (47) saturates (58) exactly at the leading order for $r\gg l_{p}$: The left-hand side becomes $\frac{\hbar\alpha_{*}}{2\pi}\frac{s_{*}}{4\pi r^{2}}-\frac{\hbar}{2\pi}\frac{1}{\sqrt{g^{*}_{rr}}}\partial_{r}\left(\frac{s_{*}}{4\pi r^{2}}\right)\approx\frac{\hbar\alpha_{*}}{2\pi}\frac{s_{*}}{4\pi r^{2}}=\frac{1}{4\pi G\eta r^{2}}$ for $r\gg l_{p}$ (from the local temperature (40) and the entropy density (42)), while the right one becomes $\langle-T^{t}{}_{t}\rangle_{*}+\langle T^{r}{}_{r}\rangle_{*}=\frac{1}{4\pi G\eta r^{2}}$ (from (IV.3)). Thus, the gravity condensate saturates not only the global bound (5) but also the local condition (55) relevant to the local structure. This is so non-trivial that there must be something behind it.
@@
The identity of black holes is still unknown, and the microscopic constituents should involve fundamental degrees of freedom in quantum gravity, where there is no notion of classical spacetime geometry. Motivated by thermodynamics and (local) holography, in the present paper, we adopted the characterization that a black hole maximizes thermodynamic entropy for a given surface area. For highly-excited spherical static configurations where the condition (18) holds, we uniquely obtained the entropy-maximizing configuration consistent with local thermodynamics and the 4D semi-classical Einstein equation with many matter fields. That is, a self-gravitating collection of near-Planckian excited quanta condensates into the radially-uniform dense configuration without horizons or singularities as in Fig.2, where the self-gravity and large quantum pressure are balanced (called semi-classical gravity condensate). The interior metric (47) (say, for $n_{s}$ scalar fields, with the parameters (33)) satisfies the semi-classical Einstein equation self-consistently and non-perturbatively in $\hbar$. The bulk dynamics makes the entropy of the interior quanta exactly follow the Bekenstein-Hawking formula, leading to the Bousso bound for thermodynamic entropy and the saturation of the local sufficient conditions. This is a candidate picture of black holes in quantum theory.
Let us discuss the ten future prospects and speculate on the possibility of a more complete quantum description of black holes.
@@
A model is obtained in the framework of group field theory, a second quantization of loop quantum gravity Oriti1 . Gluing spherically-symmetric quantum-gravity states kinematically and maximizing the entropy for a given surface area, a quantum gravitational configuration appears with the entropy proportional to the surface area, and the coefficient is fixed by using the Unruh effect and thermodynamic relations. A remarkable point is that the holographic property of the entropy of the quanta living in the interior bulk holds for any size $r$, which corresponds to the radial uniformity in our semi-classical gravity condensate. (See Sec.8 of Ref.Y1 for more discussions.)
@@
Here, interesting similarities can be observed between $n$ and $\mathcal{N}$. For our condensate, the temperature $T_{loc}^{*}(r)\sim\frac{m_{p}}{\sqrt{n}}$, characteristic wavelength $\lambda_{*}(r)\sim\sqrt{n}l_{p}$, and energy $\Delta E_{loc}^{*}(r)\sim\sqrt{n}m_{p}$ are controlled by $n$. These manifestations of $n$ are exactly the same as those of $\mathcal{N}$ in the condensate of Ref.Dvali3 , although the former is for local subsystems and the latter for the entire system. Also, from (60), the condition $n\gg 1$ corresponds to that for the classicality in Ref.Dvali3 . Therefore, a possibility is that realizing the condensate in Ref.Dvali3 in local subsystems and connecting them consistently with the semi-classical Einstein equation lead to our gravity condensate.
On the other hand, we have introduced the occupation number of excited matter quanta in the subsystem by (12): $N(r)\sim\frac{\Delta E_{loc}}{\epsilon(r)}$. For generic configurations, $N(r)\neq N_{g}(r)$ holds, while our gravity condensate satisfies
@@
Note also that the two other models above discuss quantum gravitational effects without considering quantum matter ones, while our model describes matter quanta in the self-consistent (and non-perturbative) classical gravitational field. Therefore, a more detailed study of the connection between the three models would provide some clues to the above expectation.
@@
For the gravity condensate, (62) is saturated: $\lambda_{*}(r)\sim\mathcal{R}_{*}(r)^{-1/2}$ (from (35)), and also the equivalence in the number of gravity and matter quanta (61) holds. Thus, the correspondence (63) would imply that the gravity condensate phase is governed essentially by quantum-gravitational effects, but due to the large occupation number (61) for $n\gg 1$, the mean-field approximation of quantum gravity by (1) Kiefer holds albeit barely, leading to the semi-classical description (47). This could correspond to a quantum liquid of gravity and matter.
A way to study this speculation is to first find an appropriate order parameter, study a Ginzburg-Landau-like theory or Gross-Pitaevskii-like equation including the effect of the redshift, and reproduce the radial uniformity or the interior metric. One could then consider the quantum many-body model behind it.
We now discuss a notion of global thermodynamic equilibrium in quantum gravity (related to footnote 9). In the absence of self-gravity, the most typical configurations maximize the entropy for given macroscopic parameters and correspond to the global thermodynamic equilibrium states. If this is also the case for quantum gravity, and if gravity condensates are quantum gravitational objects, then from the derivation procedure, the global thermodynamic equilibrium states in quantum gravity for a fixed surface area would be gravity condensates.
The key property in this equilibrium state is that the redshift $g_{tt}(r)\sim-e^{-\frac{R^{2}-r^{2}}{2\eta\sigma}}$ is so strong that Tolman’s law does not hold (as seen in Sec.V.1.1 and Appendix D). To understand the origin more deeply, the effect of interactions should be important. In general, a phase is a uniform equilibrium state which is achieved through internal interactions Landau_SM . Indeed, we can use a simple model and confirm that such interactions equilibrates the system in a heat bath KY2 . We could thus expect that due to many such interactions inside, energy should be transferred as conductive heat rather than radiative heat, and the propagation timescale should be consistent with $\Delta t\sim\mathcal{O}(e^{a_{0}^{2}/l_{p}^{2}})$ (see Sec.V.1.1) so that the gravity condensate exists as a phase.
@@
5. Thermodynamic entropy vs entanglement entropy—. Another field-theoretic approach to $S[g_{\mu\nu};R)$ would be to formulate it as entanglement entropy. In such cases, the Unruh temperature is often applied to local Rindler regions as the local temperature Minic ; Casini ; Jacobson_entangle , but it does not necessarily agree with the thermodynamic temperature consistent with the self-consistent gravity. In the case of, say, self-gravitating thermal radiation, the Unruh temperature is different from the local temperature obtained by applying Stefan-Boltzmann law or Tolman’s law in the metric (see Appendix A). This is natural because the local temperature of an object in global thermodynamic equilibrium, which is fixed by Tolman’s law in the self-consistent metric, does not generically coincide with the Unruh temperature determined by the acceleration required to stay against the self-gravity. Therefore, the entanglement entropy calculated by such methods can be different from the thermodynamic entropy $S[g_{\mu\nu};R)$. The crucial point is whether to take into account the self-gravity determined by the semi-classical Einstein equation.
As argued in Sec.V.1.1, however, the local temperature (40) of the gravity condensate can be obtained by applying the Unruh formula locally to the self-consistent metric (47). Therefore, we expect that applying such field-theoretic techniques locally in the metric enables us to evaluate the entropy density of the gravity condensate as entanglement entropy, and the entanglement entropy and the thermodynamic (Boltzmann) one agree and give the Bekenstein-Hawking formula. Indeed, this is the case for the quantum gravity condensate proposed in Ref.Oriti1 . We would like to check this expectation in the future.
@@
6. Physical resolution of information problem—. The purification of the initial Minkowski vacuum after evaporation and the “formal” derivation of the Page curve are only part of the information problem. More essentially, we need to understand the dynamical mechanism by which the initial wavefunction of a collapsing matter that forms a black hole is recovered after evaporation, leading to the physical Page curve as a result of the unitary evolution of matter and gravity. We also need to clarify the interior structure with quantum dynamics that resolves the singularity. Such overall consistency will reveal the true identity of black holes.
@@
In the gravity condensate, the information of a (typical) collapsing matter is stored in the bulk interior, and the amount agrees with the Bekenstein-Hawking formula. The condensate has neither horizon nor singularity (although we still need to understand the small center part at a fully quantum-gravity level), and almost all parts evaporate in the vacuum due to the Hawking-like radiation. Therefore, it should be natural to expect that (most of) the information recovers after the evaporation. As mentioned above, however, we still need to clarify the mechanism consistent with the energy flow: how the information of the matter leaks out gradually during the evaporation. A possibility is that scattering between a collapsing matter and the radiation occurs frequently at each point inside KY2 , and such interactions should transfer the information to the emitted radiation, reproducing the physical Page curve.
@@
To see this explicitly, one would at least need to check the initial-state dependence of the radiation emitted after the interaction. However, the condensate is in a typical state, and such dependence is not easy to see. One idea to overcome this is to consider perturbations from the typical state, adding a small atypical portion to the condensate and tracking it during evaporation. Another is to study a protocol that distinguishes between two typical states and implement it in this model.
@@
7. Non-typical configurations—. The gravity condensate, which is the entropy-maximizing configuration for a given surface area, can be obtained by reversible processes in a heat bath, as shown in Sec.IV.3.1. Then, what is non-typical configurations that are formed by a generic collapse of matter?
@@
One way to discuss this is to study the time evolution of a collapsing matter including the backreaction from particle creation during the collapse, considering the point in footnote 15. Indeed, we can consider the matter as consisting of many spherical shells with small energy, distributed according to the initial state, and solve (1) self-consistently KY2 . It shows that the dense structure with the metric (47) is formed only around the surface (in which (38) occurs), while the structure in deeper regions depends on the details of the initial distribution. The configuration has a smaller entropy than (43), is atypical and thus is not a black hole according to our definition.
@@
Here, it is important to consider a finite width even when discussing collapse of a spherical shell. For simplicity it is often modeled by an infinitely thin shell BD ; BHmodel , but physically, any collapsing matter is an excitation of quantum fields with a physical information $|\psi\rangle$, which due to the non-locality, cannot be localized completely in the radial direction. Considering a finite-width shell as a collection of many tiny shells and including the evaporation effect, we can see that the tangential pressure $\langle T^{\theta}{}_{\theta}\rangle$ occurs dynamically and the gravity-condensate structure appears around $r=a_{0}$, while the interior is vacuum KY2 ; KY4 . This is a result of the 4D dynamics: the 4D conservation law $\nabla_{\mu}\langle T^{\mu}{}_{\nu}\rangle=0$ contains $\langle T^{\theta}{}_{\theta}\rangle$, which can be large due to the 4D Weyl anomaly KY3 (as in Sec.IV.3.2). In the 2D cases, there is no tangential direction.
@@
Another way to answer the question is to consider a solution space, consistent with the 4D Weyl anomaly, of the semi-classical Einstein equation and study various configurations in a non-perturbative manner for $\hbar$ HKLY . It shows that the most typical ones for a fixed surface area are similar to the gravity condensate, and the number of such configurations agrees with the Bekenstein-Hawking entropy (except for $O(1)$ numerical factor). Non-typical ones have the dense structure only around the surface, which is consistent with the one above.
@@
Thus, these different studies based on the 4D non-perturbative dynamics of the semi-classical Einstein equation give almost the same picture of the typical and non-typical configurations. To clarify their relation, in the future, we would like to consider a semi-classical time-dependent perturbation and study how such non-typical ones evolve to the typical one (relaxation process) and how stable the gravity condensate is.
As an alternative approach to the stability problem, one can consider an information-theoretic version of singularity theorem: Under the assumption of the Bousso bound, singularities form when a spatial region has an entropy greater than the bound B_singularity . Therefore, if the Bousso bound is assumed and if it is possible to add some amount of entropy to the gravity condensate in a way consistent with (1), it would transform into a usual black hole with singularities, indicating an instability.
@@
8. Relation to the classical picture of black holes—. One might wonder why the gravity condensate is so different from the classical picture of black holes, a vacuum region surrounded by a horizon. The main reason for this difference is that many previous studies have not properly considered 4D quantum dynamics. First, note again that the classical one was originally derived from the classical dynamics of a collapsing matter, and that the existence of horizons has not been confirmed observationally yet. At the semi-classical level, as discussed above, infinitely thin-shell models and 2D approximation are often used to lead to almost the same picture as the classical one BHmodel . In general, however, 2D and 4D dynamics are different even in 4D spherically symmetric systems: UV divergent structures and energy conservation laws. The former provides the difference in 2D and 4D Weyl anomalies, and the latter the (non-) existence of the tangential pressure. As mentioned above, furthermore, the consideration of a finite width and the backreaction from evaporation during the collapse (consistently with footnote 15) must be considered. These points make a big difference in a self-consistent analysis of the time evolution of a collapsing matter to find a picture of black holes. Indeed, our argument KMY ; KY1 ; KY2 ; KY3 ; KY4 ; KY5 ; Y1 takes these points into account and obtains the picture of the gravity condensate.
@@
Another aspect in the difference between the conventional picture and the gravity condensate is that they belong to different branches in a solution space of the semi-classical Einstein equation. In Ref.HKLY , we constructed a self-consistent equation of the energy distribution $\frac{a(r)}{2G}$ including the effect of the 4D Weyl anomaly, and examined the structure of the solution space in a non-perturbative manner for $\hbar$. We then found that there exist two branches: one is perturbative and contains Schwarzschild-like metrics, while the other is non-perturbative and includes the dense solution (32). The point is that, the higher derivative term $\Box R_{\rm RicciScalar}$, which appears generically in the anomaly BD , causes transitions between the two branches as $r$ changes. Therefore, it would be interesting to study 4D semi-classical time evolution of a collapsing matter including such higher derivative effects and see how general the formation of the gravity condensate is and which one is formed, the conventional one or the gravity condensate.
Another interesting question is: Which is more typical, the conventional one or the gravity condensate, in the sense of greater entropy? In Sec.V.1, we have shown that the gravity condensate has the Bekenstein-Hawking entropy, including the coefficient $1/4$. Therefore, both have the same entropy and should be equally typical for a fixed surface area. To judge this in more detail, one could, say, calculate a correction term to the area entropy and find out which one is larger Y3 .
@@
9. Gravitational field with finite entropy—. We briefly comment on entropy in gravity. An interesting aspect in estimating the entropy $S[g_{\mu\nu};R)$ in Sec.II is that the local typicality and Hamiltonian constraint assign a finite entropy to a gravitational field $g_{\mu\nu}$. This is reminiscent of two things. First, it is similar to the idea that the Einstein equation corresponds to the equation of state in spacetime thermodynamics Jacobson ; Pad . Second, it should be relevant to the view of entropy as a gravitational charge Wald ; SY (see Sec.8 of Ref.Y1 for details). These could be relevant each other.